# Typical Quantum States of the Universe are Observationally Indistinguishable

### v1

@@ 3.2 Observational Indistinguishability of Typical Quantum States
for any subset $B$ of configuration space.
Now $Q_{t}$ comprises the exact positions of all particles in the universe, so observers inside a Bohmian universe cannot possibly measure $Q_{t}$; at best, they can measure a very coarse-grained version of $Q_{t}$. But the point is that even complete information about $Q_{t}$ would not help with distinguishing between typical $\Psi_{t_{0}}$ and $\rho_{0}$ (or between two typical wave functions); and a fortiori, a coarse-grained version does not help.
The essence of the difficulty with distinguishing is an instance of the general problem of deciding between two probability distributions $f_{1}$ and $f_{2}$ (say, on $\mathbb{R}^{n}$) after we are given just one point $X$ that was randomly chosen with distribution either $f_{1}$ or $f_{2}$. If $f_{1}$ and $f_{2}$ are disjoint, then this is possible; if they are nearly disjoint, then it is still possible quite reliably (for example, at the confidence level of $95\%$ often used in statistics). But if $f_{1}$ and $f_{2}$ overlap significantly, and if the observed $X$ lies in the overlap region, then this value of $X$ could have arisen from either one, so we cannot decide which $f$ was used. A reliable decision is possible only if $f_{1}(X)\approx 0$ and $f_{2}(X)$ is significantly non-zero or vice versa. Now in our case, $f_{i}=|\Psi_{i}|^{2}$ and $f_{1}\approx f_{2}$ by (5), so they overlap almost exactly and everywhere; values of $X=Q_{t}$ with $f_{1}(X)\approx 0$ will also have $f_{2}(X)\approx 0$ and therefore, with overwhelming probability, are not going to occur for either $f_{1}$ or $f_{2}$. So, the two cannot be distinguished at any time $t$ (not reliably and in fact, as we will discuss in §3.4, not even unreliably).
On the other hand, the precise microscopic trajectories of all Bohmian particles may contain more information about $\Psi_{t_{0}}$, but it is known (Dürr et al. 1992) that observers inside a Bohmian universe cannot measure positions at several times without introducing decoherence and thereby changing the trajectory. Therefore, they do not have access to the full trajectories, and the observational indistinguishability just derived remains valid. In other words, there is a known limitation to knowledge about full trajectories—the absolute uncertainty in Bohmian mechanics.
(3) OI in GRWf. We consider GRWf, the GRW theory with a flash ontology in spacetime (e.g., Tumulka 2022b, sec. 3.3.4). A key fact about OI in GRWf is this:
the joint probability distribution of all flashes in the universe up to time $t$ is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$. (8)
This follows from Theorem 1 by taking the POVM $E$ to be the POVM governing the joint distribution of all flashes (Tumulka 2022b, sec. 5.1.1 and 7.8). As in BM, observers in a GRWf universe do not have access to the full pattern of flashes, only to a coarse-grained version of it. But even the exact pattern would not provide the information needed for distinguishing a typical $\Psi_{t_{0}}$ from $\rho_{0}$ or from another typical initial wave function, for the same reasons based on Theorem 1 as discussed for BM above.
Another difference to BM is that inhabitants of a GRWf universe living at time $t$ may very well find out a lot about $\Psi_{t}$. After all, $\Psi_{t}$ is not unitarily evolved from $\Psi_{t_{0}}$ but has collapsed; for example, if we find Schrödinger’s cat alive then its wave function will have collapsed to (approximately) the wave function of a live cat. (Note that this still does not give us new information about $\Psi_{t_{0}}$.)
3.3 Remarks
Let us return to the general result of observation typicality and discuss the order of quantifiers. It has the form “for any $E$ and for most $\Psi$, $s(\Psi,E)$.” This is weaker than the statement “for most $\Psi$ and for any $E$, $s(\Psi,E)$,” which is false in this context. It is easy to confuse the two. The false but stronger statement corresponds to the following:
Super-Strong Observation Typicality
For most quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$, for any observation $E_{z}$, they cannot be distinguished by $E_{z}$.
Super-strong observation typicality implies observation typicality; it is false because if $E_{z}=|\psi\rangle\langle\psi|$ is a 1d projection in the direction of a particular $\psi\in\mathbb{S}(\mathscr{H}_{0})$, then $\mathbb{P}_{\Psi=\psi}(Z=z)=1$, whereas $\mathbb{P}_{\rho_{0}}(Z=z)=1/d_{0}\approx 0$. The super-strong property requires a uniformity in the class of typical quantum states, while observation typicality is compatible with the possibility that the typicality class depends on the choice of $E_{z}$. However, the weaker statement—observation typicality—is sufficient for our purposes. OI of typical quantum states is true for any chosen experiment, so it is true for the actual experiment that we happen to perform at an arbitrary time.
Another remark concerns alternative arguments for OI using other typicality theorems, specifically canonical typicality (Gemmer et al. 2009, Popescu et al. 2006, Goldstein et al. 2006a) and dynamical typicality as stated in (Teufel et al. 2023). We will discuss their connection to distribution typicality and observation typicality elsewhere (Chen and Tumulka 2024). Here we make two brief comments:
  1. 1.
    Suppose we have access to only local observables pertaining to a small subsystem $\mathscr{S}$ of the universe (less than half the number of degrees of freedom in the universe). Then our observations cannot distinguish typical quantum states $\Psi_{t_{0}}$ of the universe from each other. After all, canonical typicality, applied to a subspace $\mathscr{H}_{R}$ now not given by the micro-canonical energy shell (as in most discussions of canonical typicality) but by $\mathscr{H}_{R}=U_{t}\mathscr{H}_{0}$, asserts that most $\Psi_{t}\in\mathscr{H}_{R}$ have nearly the same reduced density matrix $\rho_{\mathscr{S}}=\operatorname{tr}_{\mathscr{S}^{c}}|\Psi_{t}\rangle\langle\Psi_{t}|$ in $\mathscr{S}$ (where $\mathscr{S}^{c}$ denotes the complement of $\mathscr{S}$). But the Born distribution of any observable in $\mathscr{S}$ depends on $\Psi_{t}$ only through $\rho_{\mathscr{S}}$; thus, the Born distribution accessible to observers like us is nearly independent of $\Psi_{t_{0}}$.
    In this context, the super-strong form of OI is actually true: for a typical quantum state $\Psi_{t_{0}}$, there is no local observation at all that can distinguish it from $\rho_{0}$. After all, once $\operatorname{tr}_{\mathscr{S}^{c}}|\Psi_{t}\rangle\langle\Psi_{t}|\approx\operatorname{tr}_{\mathscr{S}^{c}}(U_{t}\rho_{0}U_{t}^{\dagger})$, all observables in $\mathscr{S}$ have nearly the same Born distribution in $\Psi_{t}$ as in $U_{t}\rho_{0}U_{t}^{\dagger}$.
  2. 2.
    Dynamical typicality means here the statement that most $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$ have for every $t$ nearly the same Born distribution for all macroscopic observables (and nearly the same as $\rho_{0}$). If any scientist could distinguish $\Psi_{t_{0}}$ from $\rho_{0}$, then she or he could publish the result, say at time $t$, so there would be a macroscopic difference between $\Psi_{t_{0}}$ and $\rho_{0}$, which is impossible by dynamical typicality.
3.4 A Bayesian Analysis
So far, we have argued that no observations can reliably distinguish typical universal quantum states from each other. What about unreliable tests that may nonetheless yield probabilistic information about the quantum state? We can use the previous results to give a Bayesian analysis. Here we provide two quantitative bounds of the probabilistic information obtainable from each observation. It turns out that insofar as the observation is not too improbable, it will not give us substantial probabilistic information about the universal quantum state. Hence, there is a sense that even unreliably, universal quantum states cannot be distinguished.
@@ 3.4 A Bayesian Analysis
The assumption $d_{0}>\varepsilon^{-5}$ is reasonable; even when $\varepsilon=10^{-200}$, the condition is still satisfied as $\varepsilon^{-5}=10^{1000}\leq 10^{10^{20}}\leq d_{0}$. In this case, the posterior probability distribution differs minimally from the prior and is almost independent of any observation. This conclusion requires the assumption that the observation is not too improbable.
Hence, distribution typicality entails not only OI, but also that, in general, observations reveal very little about the initial quantum state. Notice how this fact fails in classical mechanics. Observations, even when they are probable, can reveal a great deal about the classical initial state. For example, observing the Tails outcome of a classical coin-flip experiment can rule out half the initial conditions, namely those that predict a Heads outcome. Observations cannot reveal so much in quantum mechanics because of distribution typicality and observation typicality, two facts whose analogues fail in classical statistical mechanics.
@@ 4 Comparisons with Known Results
In this section, we briefly compare our results for OI with some known results in philosophy of science and foundations of physics.
In quantum mechanics, there are several results about limitation to knowledge about the quantum state. For a survey, see (Tumulka 2022b, ch.5). Here we highlight a few:
  • Because of the Main Theorem about POVMs, we cannot measure the wave function of a given system. Relatedly, because of the no-cloning theorem, we cannot reliably copy the wave function of a given system.
  • It can be shown (Tumulka 2022a) that in any ontological theory of quantum mechanics, it is impossible to measure the ontic state. Thus, whatever the true theory of quantum mechanics may be, there must be facts that cannot be empirically determined.
  • Because of the Pusey-Barrett-Rudolph (PBR) theorem (Pusey et al. 2012), two different ensembles of wave functions with the same density matrix are physically distinct. However, they are observationally indistinguishable by all possible observations directly on the ensembles.
@@
The last item also asserts a kind of OI, but differs from our result here in that it provides a condition (equal density matrix) under which two situations (ensembles of wave functions) are indistinguishable, whereas we show here that most wave functions from $\mathscr{H}_{0}$ are indistinguishable. Moreover, all three items are compatible with substantive learning about the actual quantum state. One may still hope to rule out a significant fraction of quantum states based on some observation. Moreover, a Bayesian can, in general, update their probability distribution based on observations after which their posterior may be highly peaked. In contrast, observation typicality tells us that observations in a typical quantum universe do not rule out more than a tiny fraction of universal quantum states, and (if the observation is not too improbable) do not have a substantial influence on a uniform prior probability distribution about initial quantum states. It stays nearly uniform.
@@
We now compare our results to a much discussed underdetermination result in the philosophy of general relativity (GR). Manchak (2009) showed that for almost every spacetime model of GR, there exists another spacetime that is physically distinct (in the sense of having a different global structure) but observationally indistinguishable from it. [^7] For some philosophical discussions about this result, see Beisbart (2009), Norton (2011), Butterfield (2012), and Cinti and Fano (2021).
Both Manchak’s result and ours are compatible with a number of inductive hypotheses. In Manchak’s case, one can assume that the same law of nature (Einstein’s equation of GR) applies to all spacetime points. Moreover, Manchak’s result is compatible with any set of local conditions (conserved under local isometries). In our case, one can also assume that the same laws of nature (within quantum theory) apply to the entire model, so that we are considering the same Hilbert space $\mathscr{H}$, the same Hamiltonian $H$, and the same subspace $\mathscr{H}_{0}$.
There are important differences too. First, Manchak’s argument shows the existence of an OI counterpart for almost every model of GR, while our argument shows that almost every model is an OI counterpart to almost every other model of QM. Hence, his result is compatible with the fact that we can observationally rule out a significant fraction of possible universes while ours is not. Second, Manchak’s proof uses a “cut-and-paste” method to construct special models of GR that are OI from the original model, while ours uses the probabilistic method to show that OI must hold among typical models of QM with generic features.
@@
Related to the second point, Manchak’s argument has been challenged on philosophical grounds that the constructed spacetimes are not “physically reasonable,” because some people regard the “cut-and-paste” models as implausible candidates for the actual universe. (For example, see Cinti and Fano (2021).) One may respond, as Manchak does, that “the theorem can have physical relevance even if the particular model constructed in the proof does not” (Manchak 2011, p.7). We do not take a stance on that issue, but merely point out that the objection of “physical unreasonableness” does not even work as a prima facie objection against our argument for OI in QM. We do not show that almost any model has at least one specially constructed counterpart. Rather, we show that typical (generic) models of QM in a high-dimensional Hilbert space are OI from each other.
It is interesting to note that GR and QM, two pillars of modern physics, both lead to OI.

### v2

@@ 3.2.2 Not Even Unreliably Distinguishable
Then, since the number of outcomes would realistically be less than $10^{10^{30}}$, the probabilities less than $\varepsilon$ would add up to less than $10^{-9\cdot 10^{30}}$. Thus, outcomes with probability less than $\varepsilon$ will not be observed given the computational limitations, and still our examples satisfy $d_{0}=10^{10^{80}}\gg 10^{3\cdot 10^{31}}=\varepsilon^{-3}$.
3.2.3 A Bayesian Analysis
Theorem 1 also allows us to perform a Bayesian analysis, i.e., to mathematically express the probabilistic information provided by the observation of the outcome $z$. Here, we provide two quantitative bounds on this information. They have the consequence that, if the observation is not too improbable, it cannot yield substantial probabilistic information about the universal quantum state. This result invites further analysis of its implications for Bayesian epistemology and convergence to the truth (see Lin (2024) for a survey), which we leave for future work.
@@ 3.2.3 A Bayesian Analysis
The assumption $d_{0}>\varepsilon^{-5}$ is mild; for example for (12), $\varepsilon^{-5}=10^{5\cdot 10^{31}}$ is far less than (5).
The physical consequence of these corollaries will be that, in general, observations reveal very little about the initial quantum state.
Example 4.
To illustrate that the assumption of large $d_{0}$ cannot be dropped, we compute $f(\psi)$ explicitly in a case with $d_{0}=2$: Let $\{|b_{1}\rangle,|b_{2}\rangle\}$ be an orthonormal basis of $\mathscr{H}_{0}$, and let $E_{z}=|b_{1}\rangle\langle b_{1}|$. Then [^15] Proof: $\mathbb{P}(Z=z|\Psi=\psi)=\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}$ by Born’s rule, so $\mathbb{P}(Z=z,\Psi\in d\psi)=\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}u_{0}(d\psi)$ and $\mathbb{P}(Z=z)=1/2$ by symmetry, so $\mathbb{P}(\Psi\in d\psi|Z=z)=2\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}u_{0}(d\psi)=f(\psi)\,u_{0}(d\psi)$, which yields (15).
$f(\psi)=2\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}\,,$ (15)
which varies between 0 and 2 and deviates significantly from 1 for a substantial fraction of $\psi$’s: indeed, it follows from Example 2 that $f(\psi)$ exceeds 1.8 for 10% of all $\psi$’s.
Example 5.
To illustrate that the assumption $\operatorname{tr}(\rho_{0}E_{z})>\varepsilon$ cannot be dropped, we consider an orthonormal basis $\{|b_{1}\rangle,\ldots,|b_{d_{0}}\rangle\}$ of $\mathscr{H}_{0}$ and the POVM $E_{z}=|b_{z}\rangle\langle b_{z}|$ for $z=1,\ldots,d_{0}$. Then $\operatorname{tr}(\rho_{0}E_{z})=1/d_{0}$ for any $z$, which is much less than $\varepsilon$ if $d_{0}>\varepsilon^{-5}$. By the same reasoning as for (15), $f(\psi)=d_{0}|\langle\psi|b_{z}\rangle|^{2}$, which is far from being constantly 1.161616Proof: It is known that for random $\Psi$ with uniform distribution and large $d_{0}$, $\langle\Psi|b_{z}\rangle$ has an approximately Gaussian distribution in the complex plane with mean 0 and variance $\mathbb{E}[|\langle\Psi|b_{z}\rangle|^{2}]=1/d_{0}$. Since this distribution gives measure 1/4 to the disk around the origin with radius $\sqrt{\ln(4/3)/d_{0}}$ and 1/4 to the complement of the disk around the origin with radius $\sqrt{\ln(4)/d_{0}}$, we have that $f(\psi)<\ln(4/3)<0.288$ for 1/4 of all $\psi$’s and $f(\psi)>\ln(4)>1.38$ for 1/4 of all $\psi$’s.
3.3 Physical Consequences: First Argument for OI
We now explain how these mathematical facts imply OI in the sense of observation typicality in OQM, BM, GRW theory, and EQM. We have two arguments, the first of which will be presented now, the second in §3.4; a comparison of the two will be made in §3.5.1.
The first argument is based on the Main Theorem about POVMs (see §2.1), which is known to be valid in OQM, BM, GRW theory, and EQM. The argument puts together the steps we have laid out already. By the Past Hypothesis, $\Psi_{t_{0}}$ is typical in $\mathscr{H}_{0}$, which has dimension $d_{0}\geq 10^{10^{80}}\gg\varepsilon^{-5}$ even for $\varepsilon=10^{-10^{31}}$. Suppose we carry out any experiment $\mathscr{E}$; $\mathscr{E}$ is associated, according to the Main Theorem about POVMs, with some POVM $E$; suppose we obtained the outcome $Z=z$ associated with the operator $E_{z}$. Then according to Corollary 1, typical quantum states in $\mathscr{H}_{0}$ are not reliably distinguished by this outcome from $\rho_{0}$. Suppose further $\operatorname{tr}(\rho_{0}E_{z})>\varepsilon$. By Corollary 2, the outcome does not significantly favor one $\Psi$ over another or over $\rho_{0}$. And by Corollaries 3 and 4, the Bayesian posterior distribution of $\Psi$ does not significantly differ from the prior. Thus, regardless of which experiment observers are making, the outcome $z$ provides no significant information about $\Psi$. This completes the argument.
This reasoning uses only one outcome $z$ and the associated operator $E_{z}$; it does not matter how many other $z^{\prime}$ there are and what the $E_{z^{\prime}}$ are. OI of typical quantum states is true for any experiment, so it is true for the actual experiment that we happen to perform at an arbitrary time with $z$ standing in for the actual outcome.
In quantum theory with unitary evolution, OI extends to later states too: Since typical $\Psi_{t}$ from the appropriately evolved subspace $U_{t}\mathscr{H}_{0}$ lead to nearly identical probability for outcome $Z=z$, the observation of this outcome does not distinguish among different $\Psi_{t}$.
3.4 Physical Consequences: Second Argument for OI
Our second argument also works in all of OQM, BM, GRW theory and EQM. It is more direct than the first argument, particularly in BM and the flash version of GRW. We will formulate it first in BM, and afterwards comment on its translation to the other versions of QM.
The argument concerns the configuration $Q_{t}$ of the universe at an arbitrary time $t$; $Q_{t}$ is typical with respect to the $|\Psi_{t}|^{2}$ distribution, and this distribution is given in terms of the initial wave function $\Psi_{t_{0}}$ of the universe by a POVM (in fact, a PVM (projection-valued measure)) $P$,
[…]
for any subset $\Omega$ of the configuration space of the universe. This POVM $P$ is the Heisenberg-evolved configuration observable, i.e., the projection-valued measure jointly diagonalizing all position operators at time $t$, or $P(\Omega)=U_{t}^{\dagger}1_{\Omega}U_{t}$, where $1_{\Omega}$ is the multiplication operator multiplying by the characteristic function of $\Omega$.
Now $Q_{t}$ comprises the exact positions of all particles in the universe, so observers inside a Bohmian universe cannot possibly measure $Q_{t}$; at best, they can measure a very coarse-grained version of some components of $Q_{t}$. But the crucial point here is that any information available to the inhabitants of a Bohmian universe at time $t$ will be recorded in $Q_{t}$, in fact in whether $Q_{t}$ lies in some (suitably coarse, macroscopic) set $\Omega$. For example, observations (and conclusions drawn from them) will be reported in books and journals, and thus in the configuration of ink particles. (For more detailed discussions, see (Albert 2015, pp. 83-85).) More precisely, we make use of the following
Necessary Condition
If an event is known to inhabitants of a Bohmian universe at time $t$ then it is of the form $Q_{t}\in\Omega$ for some set $\Omega$.
The argument for OI in BM consists of applying Corollaries 14 to the POVM $\{E_{A},E_{B}\}=\{P(\Omega),P(\Omega^{c})\}$ for some subset $\Omega$ of the configuration space $\mathbb{R}^{3N}$ of the universe and its complement $\Omega^{c}=\mathbb{R}^{3N}\setminus\Omega$. As a consequence of the necessary condition just mentioned and Corollaries 14, no event known to the inhabitants can reliably distinguish $\Psi_{A}$ from $\rho_{0}$, or can significantly favor $\Psi_{A}$ over $\rho_{0}$ or $\Psi_{B}$, or can substantially alter the Bayesian posterior distribution of $\Psi_{t_{0}}$ (for the last two disjuncts provided $\operatorname{tr}(\rho_{0}P(\Omega))>\varepsilon$, i.e., the event $Q_{t}\in\Omega$ is not too improbable on average). Thus, at no time $t$ can inhabitants of a Bohmian universe have significantly more information about $\Psi_{t_{0}}$ than what is already provided by the Past Hypothesis. This completes the argument.
We remark that the precise microscopic trajectories of all Bohmian particles may contain more information about $\Psi_{t_{0}}$, but it is known that observers inside a Bohmian universe cannot measure positions at several times without introducing decoherence and thereby changing the trajectory. Therefore, they do not have access to the full trajectories, which makes understandable why the dependence of the trajectories on $\Psi_{t_{0}}$ is compatible with OI as just derived.
Let us summarize the second argument in a different way. A key fact in BM that leads to OI is this:
[…]
Corollaries 14 just provide quantitative statements (in the more general framework of POVMs) of how this near-independence makes $\Psi$ observationally indistinguishable from $\rho_{0}$ or other $\Psi$s. They show that that even complete information about $Q_{t}$ would not help with distinguishing between typical $\Psi_{t_{0}}$ and $\rho_{0}$ (or between two typical wave functions); a fortiori, limited information about $Q_{t}$ does not help.
In OQM, the second argument can be applied as well (although it is a bit less clear because the ontology of OQM is less clear): after all, the Heisenberg cut (between system and apparatus) can always be moved so as to enlarge the system, and the $|\Psi_{t}|^{2}$ distribution in configuration space will also in OQM determine the distribution of macroscopic records (say, in books or journals) at time $t$. Thus, the macroscopic records at $t$ provide almost no information about $\Psi_{t_{0}}$.
Concerning EQM, we have reservations about the justification of the use of probabilities (Barrett 2023, sec. 6; Tumulka 2022b, sec. 3.5.5); but to the extent that they can be justified, our second argument for OI applies as well.
In GRW theory, the second argument can be applied in a slightly modified variant. To this end, we consider GRWf, the GRW theory with a flash ontology in spacetime (Ghirardi and Bassi 2020, sec. 11; Tumulka 2022b, sec. 3.3.4). Let $F_{t}$ denote the pattern of flashes up to $t$, with $t>t_{0}$. A key fact about OI in GRWf is this:
for any subset $M\in\cup_{n=0}^{\infty}\mathbb{R}^{4n}$, the probability of the event that $F_{t}\in M$ is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$. (18)
This follows from Theorem 1 by taking the POVM $E$ to be the POVM governing the joint distribution of all flashes (Tumulka 2022b, sec. 5.1.1 and 7.8). As for the configuration in BM, observers in a GRWf universe do not have access to the full pattern of flashes, $F$, at most to a coarse-grained version of it. But again, we use the fact that any record would have to be encoded in the flashes (as are the positions of pointers or the shape of printed letters). More precisely, we use the following
Necessary Condition
If an event is known to inhabitants of a GRWf universe at time $t$ then it is of the form $F_{t}\in M$ for some set $M$.
Now Corollaries 14 show that no such event can reliably distinguish $\Psi_{A}$ from $\rho_{0}$, or can significantly favor $\Psi_{A}$ over $\rho_{0}$ or $\Psi_{B}$, or can substantially alter the Bayesian posterior distribution of $\Psi_{t_{0}}$, provided $\operatorname{tr}(\rho_{0}E(M))>\varepsilon$. Thus, at no time $t$ can observers in a GRWf universe have significant information about $\Psi_{t_{0}}$. This completes the argument.
Another difference to BM is that inhabitants of a GRWf universe living at time $t$ may very well find out a lot about the wave function $\Psi_{t}$ at their time (in contrast to the initial wave function $\Psi_{t_{0}}$). After all, $\Psi_{t}$ is not unitarily evolved from $\Psi_{t_{0}}$ but has collapsed; for example, if we find Schrödinger’s cat alive then its wave function will have collapsed to (approximately) the wave function of a live cat. (Note that this still does not give us new information about $\Psi_{t_{0}}$.)
3.5 Remarks
3.5.1 Experiment Framing vs $|\Psi_{t}|^{2}$ Framing
Our first remark concerns a comparison of the two arguments given in §3.3 and §3.4. Although both used the Corollaries 14, they used them in different ways: the first was framed in terms of experiments, the second in terms of the $|\Psi_{t}|^{2}$ distribution in configuration space. The first argument pretends that the observers are somehow not part of the universe. Although we do not think that this approach leads to any wrong conclusions here, we note that the second argument is clear-cut and the first less so, as various questions can arise about the first: Is it not usually assumed that observers and their apparatus are not entangled with the object they are observing? And is that not violated in the relevant observations of the universe? Given that in OQM, observations can collapse wave functions, should not $E_{z}$ be applied to a collapsed wave function instead of $\Psi_{t_{0}}$? While none of these objections to the first argument withstands scrutiny, the second argument makes it much more transparent that the conclusion, OI, remains valid. The second argument also makes it clear that repeating the experiment does not allow the observer to measure the probability more precisely; after all, records of all experiments are included in the configuration at time $t$, and the whole universe cannot be repeated, of course.
3.5.2 Other Kinds of Indistinguishability
As already remarked, the question we ask is different from the question, given two states $\Psi_{A},\Psi_{B}$, is there a POVM that will distinguish them? It is also different from the question, can we empirically distinguish the possibility that the universe is in the mixed state $\rho_{0}$ from the possibility that it is in a random pure state $\Psi$ with distribution $u_{0}$? The answer to the latter question is trivially “no” for any value of $d_{0}$ (large or not), as any observation $z$ with POVM element $E_{z}$ has probability $\operatorname{tr}(\rho_{0}E_{z})$ in the mixed case and probability $\int_{\mathbb{S}(\mathscr{H}_{0})}u_{0}(d\psi)\langle\psi|E_{z}|\psi\rangle$ in the random pure case, and the two probabilities are always exactly equal. (Readers may even find it hard to see any physical difference at all between the mixed state $\rho_{0}$ and a random pure state $\Psi$ with distribution $u_{0}$; but this difference becomes visible in Bohmian mechanics, where trajectories guided by $\rho_{0}$ are different from those guided by $\Psi$.)
Our results convey that the Born distribution $\langle\Psi|E_{z}|\Psi\rangle$ is close to $\operatorname{tr}(\rho_{0}E_{z})$ for most $\Psi$, whereas the impossibility to distinguish the mixed case from the random pure case arises from the fact that $\langle\Psi|E_{z}|\Psi\rangle$, when averaged over $\Psi$, yields $\operatorname{tr}(\rho_{0}E_{z})$. The difference between our results and the impossibility to distinguish the mixed case from the random pure case is perhaps most easily visible in the Bayesian statement of §3.2.3.
@@ 4 Comparisons with Known Results
We compare observation typicality with some known results in foundations of physics.
There are several results about limitation to knowledge about the quantum state. For a recent survey, see (Tumulka 2022b, ch.5). Here we highlight a few: (1) Because of the Main Theorem about POVMs, we cannot measure the wave function of a given system with useful precision. Relatedly, because of the no-cloning theorem, we cannot reliably copy the wave function of a given system. (2) It can be shown (Tumulka 2022a) that in any ontological theory of quantum mechanics, it is impossible to measure the ontic state. Thus, whatever the true theory of quantum mechanics may be, there must be facts that cannot be empirically determined. (3) Because of the Pusey-Barrett-Rudolph (PBR) theorem (Pusey et al. 2012), two different ensembles of wave functions with the same density matrix are physically distinct. However, they are observationally indistinguishable by all possible observations directly on the ensembles.
@@
The last item also asserts a kind of OI, but differs from our result here in that it provides a condition (equal density matrix) under which two situations (ensembles of wave functions) are indistinguishable, whereas we show here that most wave functions from $\mathscr{H}_{0}$ are indistinguishable. Moreover, all three items are compatible with substantive learning about the actual quantum state. One may still have hoped to reliably eliminate a significant fraction of quantum states based on some observation. One might also have hoped that, a Bayesian could, in general, update their probability distribution based on observations, leading to a highly peaked posterior. However, observation typicality conveys that these are impossible for the universal quantum state. Observations in a typical quantum universe do not reliably eliminate more than a tiny fraction of universal quantum states, and—if the observation is not too improbable—do not significantly reshape a uniform prior probability distribution over initial quantum states. The distribution remains nearly uniform.
@@
In the philosophy of general relativity (GR), Manchak (2009, 2011) showed that for every spacetime model of GR of a certain class, there exists another spacetime that is physically distinct (in the sense of having a different global structure) but observationally indistinguishable from it. [^17] For some philosophical discussions about this result, see Beisbart (2009), Norton (2011), Butterfield (2012), and Cinti and Fano (2021). Remarkably, both GR and QM, the two pillars of modern physics, imply OI.
@@
We highlight two differences between the arguments. First, Manchak’s argument shows the existence of an OI counterpart for every model of GR in a special class. Hence, his result is compatible with the fact that we can observationally rule out a significant fraction of possible universes while ours is not. Second, Manchak’s proof uses a “cut-and-paste” method to construct special models of GR that are OI from the original model, while ours uses the probabilistic method to show that OI must hold among typical models of QM with generic features. Related to this point, Manchak’s argument has been challenged on philosophical grounds that the constructed spacetimes are not “physically reasonable,” because some people regard the “cut-and-paste” models as implausible candidates for the actual universe. (For example, see Cinti and Fano (2021).) We do not take a stance on that issue, but merely point out that the objection of “physical unreasonableness” does not even work as a prima facie objection against our argument for OI in QM. We do not show that almost any model has at least one specially constructed counterpart. Rather, we show that typical (generic) models of QM in a high-dimensional Hilbert space are OI from each other.